OpenStudy (anonymous):

Jose opens a savings account with a $100 deposit. He then deposits $8.00 each week. In how many weeks will he have a total of $2,000?

OpenStudy (anonymous):

2000 = 100 + 8x 1900 = 8x x = 1900/8 = 237.5 Since you jose deposits every week you must round up. He would have to deposit 8 dollars every week for 238 weeks.

OpenStudy (anonymous):

Thank you!